OpenStudy (anonymous):

Can someone help me solve this 1/x +5=7

pooja195 (pooja195):

Multiply by the reciporcal this will help cancel the fraction \[\huge~\rm~\bf~1/x(x)+5(x)=7(x)\]

pooja195 (pooja195):

What do we get once we do this? Right the next equation

OpenStudy (anonymous):

Do you mean to multiply by x to 7 too

pooja195 (pooja195):

Yes

pooja195 (pooja195):

Did you figure out the next line? :)

OpenStudy (mathmale):

I propose that it'd be a lot easier to combine the 5 and 7 first. If 1/x +5=7, subtract 5 from both sides of the equation. You get ... what?

pooja195 (pooja195):

Either way you get the same response. I just did the harder part first.

OpenStudy (mathmale):

Let's let @thevi respond, please.

pooja195 (pooja195):

I think we confused them.

jaynator495 (jaynator495):

HEY HEY MATHMALE! HEY HEY! HOW ARE YOU LONG TIME NO SPEAK! XD

OpenStudy (mathmale):

Never mind. @thevi is now offline.

pooja195 (pooja195):

@Jaynator495 refrain from spamming posts.

TheSmartOne (thesmartone):

The asker already asked this question before, and I *attempted* to help them, but they never responded.

OpenStudy (anonymous):

Mat male thanks I got the answer your way. But thank you both of you for helping me

OpenStudy (mathmale):

You're welcome.

pooja195 (pooja195):

:)

TheSmartOne (thesmartone):

I believe the asker is using the app, that's why it shows them as offline.

pooja195 (pooja195):

What did you get as the answer?

OpenStudy (anonymous):

Thesmartone sorry I thought that question didn't get posted so I posted another one I gave u a best response Thanks

TheSmartOne (thesmartone):

Oh, it's ok :)